OpenStudy (anonymous):

Suppose the population in a town has an average growth rate of 18%. If there are 15,500 people now living in town in about how many years will the population grow to 30,000? Use the equation 30,000=(15,500)(1.18)^x and round the value of x to the nearest year

OpenStudy (cwrw238):

This is an equation with the variable as an exponent. Have you come across this type before?

OpenStudy (cwrw238):

first isolate the exponential bit by dividing both sides by 15,500 (1.18)^x = 30,000/15,500

OpenStudy (cwrw238):

now you take logarithms of both sides ln (1.18)^x = ln (30,000/15500) can you continue from here? Hint;- use the rule ln a^b = b ln a